# Universal Hitchin moduli spaces

### v1

@@ 5.1. The universal Higgs field
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$. Consider $h\in\Omega^{0}(E_{G}(G/K))$ a reduction of structure group of $E_{G}$ to $K$, and let $E_{K}$ be the corresponding principal $K$-bundle. Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation. Consider the space
$\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})).$
@@
$\curly{X}^{Higgs}=\{(J,A,\psi)\in\curly{X}\;|\;\bar{\partial}_{J,A}\varphi=0\}$
@@
for $(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$. Notice that this complex structure corresponds to the natural fibrewise cotangent complex structure in the holomorphic fibration $\curly{J}\times{\curly A}\to\curly{J}$. In particular, the map $\pi_{1}\colon\curly{X}\to\curly{J}$ is holomorphic, and the complex structure along the fibres coincides with the $J$-dependent complex structure $\mathbf{I}$ in (2.8). Using the Chern-Singer correspondence, it is not difficult to give holomorphic coordinates on the space $\curly{J}\times{\curly A}\cong\curly{J}\times{\curly C}$, and hence on $(\curly{X},\mathbb{I})$. In the next result we give an independent proof of the vanishing of the Nijenhuis tensor $N_{\mathbb{I}}$ of $\mathbb{I}$.
@@
The complex structure $\mathbb{I}$ is formally integrable, that is, its Nijenhuis tensor vanishes
@@
The complex structures on the fibers and base of the fibration $(\curly{X},\mathbb{I})\to(\curly{J},\mathbb{J}{J})$ are well-known to be integrable. Hence, it suffices to calculate the mixed component of the Nijenhuis tensor
@@
$=(0,-a(J\dot{J}+\dot{J}J),\dot{\psi}(J\dot{J}+\dot{J}J))=0.$
In our next result we show that $\curly{X}^{Higgs}\subset(\curly{X},\mathbb{I})$ is fomally a complex submanifold.
@@
The tangent space of $\curly{X}^{Higgs}$ at $(J,A,\psi)$, is given by triples $(\dot{J},a,\dot{\psi})\in T_{J}\curly{J}\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(X,E_{K}(\mathfrak{k}))$ satisfying
@@
Consequently, $\curly{X}^{Higgs}$ is fomally a complex submanifold of $(\curly{X},\mathbb{I})$.
@@
$=i\left(d_{A}\dot{\psi}^{1,0_{J}}+[a,\varphi]-\frac{i}{2}d_{A}\left(\psi(\dot{J})\right)\right)=0$
@@
Our next goal is to define a universal moduli space of Higgs bundles, with a holomorphic fibration structure over Teichmüller space. The first step is to show that the integrability condition for $\varphi$ is compatible with the group $\widetilde{\curly{G}}$ defined in (4.1). We observe first that there is a holomorhic action of $\widetilde{\curly{G}}$ on $\curly{J}\times{\curly A}$, given by
@@
where $A_{g}$ is the defined via the Chern-Singer correspondence
@@
$\curly{X}^{ps}\subset\curly{X}^{Higgs}\subset\curly{X}$
@@
By construction, $\mathcal{U}^{Higgs}(G)$ fibers over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
(5.3) $\mathcal{U}^{Higgs}(G)\to\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
@@ 5.2. Kähler fibration for the universal Higgs field
Consider the structure of Kähler fibration $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ on $(\curly{X},\mathbb{I})\to\curly{J}$ defined fibrewise by the symplectic structure $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ on ${\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$. More explicitly,
@@
The aim of this section is to prove that such structure admits a Kähler Ehresmann connection compatible with Hitchin’s equations (2.15). The situation here is opposite to that on Section 3.2: the holomorphic fibration defined by $\mathbb{I}$ is non-trivial, whereas $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ defines a trivial symplectic fibration, constant along $\curly{X}$. Nonetheless, the combination of these two structures defines a non-trivial structure of Kähler fibration on $\curly{X}\to\curly{J}$, different from the one considered in Section 3.2.
In order to achieve our goal in this section, we will apply Theorem 3.2. As we have seen in Lemma 5.2, the integrability condition
@@
cuts a holomorphic submanifold $\curly{X}^{Higgs}$ of $(\curly{X},\mathbb{I})$. Given that $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ is constant on $\curly{X}$, it provides a candidate for coupling two-form $\boldsymbol{\sigma}$ on $(\curly{X},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$, inducing the trivial Ehresmann connection on $\curly{X}\to\curly{J}$, regarded as the product $\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k})$. However, one can readily check that the horizontal subspace of this connection is not tangent to the submanifold $\curly{X}^{Higgs}$, and hence is not appropiate for our goals. Motivated by this, we consider the Ehresmann connection $\Gamma^{\mathbb{I}}$ with horizontal subspace
$H^{\mathbb{I}}=\Big\{\left(\dot{J},0,-\tfrac{1}{2}\psi(J\dot{J})\right)\;|\;\dot{J}\in T_{J}\curly{J}\Big\}.$
@@
where $\mathbf{I}$ denotes the complex structure along the fibres. Furthermore, its curvature $F_{\mathbb{I}}\in\Omega^{2}(\curly{X},V\curly{X})$ is of type $(1,1)$ and given explicitly by
@@
(5.6) $\begin{split}\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(\Gamma^{\mathbb{I}}v_{1},\Gamma^{\mathbb{I}}v_{2})&=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{1}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{1}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right)\end{split}$
Our next result proves that the Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler, by calculation of an explicit coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ via equation (3.3). Define a basic two-from $\mu\in\Omega^{2}(\curly{X})$ by
@@
The Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler. Furthermore, the following formula gives a coupling form for $\Gamma^{\mathbb{I}}$
@@
$=-\widehat{\boldsymbol{\omega}}_{\mathbf{I}}((0,a,\dot{\psi}),F_{\mathbb{I}}(v_{1},v_{2})).$
To finish this section, we provide a formula for the the symmetric tensor on $\curly{X}$ associated to the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$, explicitly given by
@@
By construction, this coincides with the flat hyperkähler metric (2.11) along the fibres of $\curly{X}\to\curly{J}$. Similarly as in Corollary 3.7, $\mathbf{g}_{\mathbb{I}}$ is negative semi-definite along the horizontal directions of the connection $\Gamma^{\mathbb{I}}$.
@@
For any tangent vector $v=(\dot{J},a,\dot{\psi}),\in T_{(J,A,\psi)}\curly{X}$ one has
(5.8) $\begin{split}\mathbf{g}_{\mathbb{I}}(v,v)&=\int_{\Sigma}B(a\wedge Ja)+\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{1}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{1}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right)\end{split}$
@@
Formula (5.8) follows from (5.7), using that $\Gamma^{\mathbb{I}}\circ\mathbb{I}=\mathbf{I}\circ\Gamma^{\mathbb{I}}$ (see Lemma 5.4).
5.3. Hamiltonian action and coupled Hitchin’s equations
We follow the notation of the previous section. Our aim in this section is to investigate a natural Hamiltonian action for minimal couplings associated to the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ for the connection $\Gamma^{\mathbb{I}}$, in the Kähler fibration $(\curly{X},\mathbb{I},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$ over the space of complex structure $\curly{J}$.
We fix a symplectic form $\omega$ on $\Sigma$, compatible with the orientation. Given $\alpha>0$ a real coupling constant and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
@@ 5.3. Hamiltonian action and coupled Hitchin’s equations
The following result is the analogue of Lemma 3.8 in the present setup. The proof is analogue, by application in this case of Corollary 5.7, and is ommitted.
@@
Given a horizontal vector field $v\in H^{\mathbb{I}}$ at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$, one has
@@
    For $\varepsilon=-1$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is negative definite along $H^{\mathbb{I}}$. In particular, $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,-1}$ is a non-degenerate symplectic structure.
@@
    For $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
Our next goal is to prove that $(\curly{X},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ admits a Hamiltonian action by the extended gauge group $\widetilde{{\curly K}}$, determined by $\omega$ and the reduction $E_{K}\subset E_{G}$. For this, we will need some refined information about the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$. Consider the closed two form $\pi^{*}_{2}\omega_{\mathbf{I}}$ on $\curly{X}$, given by the pull-back of $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ via the canonical projection $\pi_{2}\colon\curly{X}\to{\curly A}\times\Omega^{1}(X,E_{K}(\mathfrak{k}))$. Explicitly, for $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j})\in T_{(J,A,\psi)}\curly{X}$, $j=1,2$, one has
@@
The two-form $\pi^{*}_{2}\omega_{\mathbf{I}}$ restricts on the fibres to $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$, and hence, by the general theory of symplectic fibrations (see [GLS, Section 1.6]), one expects that there exists a one-form $\lambda\in\Omega^{1}(\curly{X})$ such that
@@
Consider the one-form $\lambda\in\Omega^{1}(\curly{X})$ defined by
@@
for $v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$. Then, formula (5.13) holds. Furthemore, considering the $L^{2}$-norm of the unitary Higgs field
@@
(5.14) $\boldsymbol{\sigma}^{\mathbb{I}}=\pi^{*}_{2}\omega_{{\curly A}}+\tfrac{i}{2}\partial_{\mathbb{I}}\bar{\partial}_{\mathbb{I}}\nu$
@@
$=-4\pi^{*}_{2}\omega_{{\curly A}}+4\boldsymbol{\sigma}^{\mathbb{I}}.$
Consider the extended gauge group $\widetilde{{\curly K}}$ of $h$ and $\omega$. Similarly as in [AGG, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{I},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ preserving $\mathbb{I}$ and $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ by push-forward. We are ready to prove the main result of this section. We use the same notation as in Proposition 3.9.
@@
$-\int_{\Sigma}f\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)$
@@
$=-\int_{\Sigma}fd^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)$
Similarly as in Proposition 3.9, the $\widetilde{{\curly K}}-$action produces a coupling term in the base $\Sigma$ (cf. Remark 4.2), which combines with the scalar curvature of the metric $g=\omega(,J)$. In particular, similarly as in [AGG], zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{I}}$ are given by solutions of the coupled system of equations
(5.15) $\begin{split}F_{A}-\tfrac{1}{2}[\psi\wedge\psi]&=0\\ \varepsilon S_{J}+\alpha\Lambda_{\omega}\left(d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V},\end{split}$
In the next section we combine these equations with the integrability condition $\bar{\partial}_{A}\varphi=0$ for the Higgs field, in order to introduce the universal moduli space of Hitchin’s equations.

### v2

@@ 5.1. The universal Higgs field
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$, $h\in\Omega^{0}(E_{G}(G/K))$ a reduction of the structure group of $E_{G}$ to $K$, and $E_{K}$ the corresponding principal $K$-bundle. Let $\curly{J}$ be the space of complex structures on $\Sigma$ compatible with the given orientation. Consider the space
$\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k})).$
@@
$\curly{X}^{Higgs}=\{(J,A,\psi)\in\curly{X}\;|\;\bar{\partial}_{J,A}\varphi=0\},$
@@
for $(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$. Notice that this complex structure corresponds to the natural fibrewise cotangent complex structure in the holomorphic fibration $\curly{J}\times{\curly A}\to\curly{J}$. In particular, the map $\pi_{1}\colon\curly{X}\to\curly{J}$ is holomorphic, and the complex structure along the fibres coincides with the $J$-dependent complex structure $\mathbf{I}$ in (2.7). Using the Chern–Singer correspondence, it is not difficult to give holomorphic coordinates on the space $\curly{J}\times{\curly A}\cong\curly{J}\times{\curly C}$, and hence on $(\curly{X},\mathbb{I})$. In the next result we give an independent proof of the vanishing of the Nijenhuis tensor $N_{\mathbb{I}}$ of $\mathbb{I}$.
@@
The complex structure $\mathbb{I}$ is formally integrable, that is, its Nijenhuis tensor vanishes:
@@
The complex structures on the fibres and base of the fibration $(\curly{X},\mathbb{I})\to(\curly{J},\mathbb{J}{J})$ are well-known to be integrable. Hence, it suffices to calculate the mixed component of the Nijenhuis tensor
@@
$=(0,-a(J\dot{J}+\dot{J}J),\dot{\psi}(J\dot{J}+\dot{J}J))=0.\qed$
In our next result we show that $\curly{X}^{Higgs}\subset(\curly{X},\mathbb{I})$ is formally a complex submanifold.
@@
The tangent space of $\curly{X}^{Higgs}$ at $(J,A,\psi)$, is given by triples $(\dot{J},a,\dot{\psi})\in T_{J}\curly{J}\oplus\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))\oplus\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$ satisfying
@@
Consequently, $\curly{X}^{Higgs}$ is formally a complex submanifold of $(\curly{X},\mathbb{I})$.
@@
$=i\left(d_{A}\dot{\psi}^{1,0_{J}}+[a,\varphi]-\frac{i}{2}d_{A}\left(\psi(\dot{J})\right)\right)=0,$
@@
Our next goal is to define a universal moduli space of Higgs bundles, with a holomorphic fibration structure over Teichmüller space. The first step is to show that the integrability condition for $\varphi$ is compatible with the group $\widetilde{\curly{G}}$ defined in (4.1). We observe first that there is a holomorphic action of $\widetilde{\curly{G}}$ on $\curly{J}\times{\curly A}$, given by
@@
where $A_{g}$ is the defined via the Chern–Singer correspondence:
@@
$\curly{X}^{ps}\subset\curly{X}^{Higgs}\subset\curly{X},$
@@
By construction, $\mathcal{U}^{Higgs}(G)$ fibres over the Teichmüller space $\mathcal{T}$ of complex structures modulo diffeomorphisms isotopic to the identity
(5.3) $\mathcal{U}^{Higgs}(G)\longrightarrow\mathcal{T}:=\curly{J}/\operatorname{Diff}_{0}(\Sigma).$
@@ 5.2. Kähler fibration for the universal Higgs field
Consider the structure of Kähler fibration $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ on $(\curly{X},\mathbb{I})\to\curly{J}$ defined fibrewise by the symplectic structure $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ on ${\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$. More explicitly,
@@
The aim of this section is to show that such a structure admits a Kähler Ehresmann connection compatible with Hitchin’s equations (2.12). The situation here is opposite to that in Section 3.2: the holomorphic fibration defined by $\mathbb{I}$ is non-trivial, whereas $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ defines a trivial symplectic fibration, constant along $\curly{X}$. Nonetheless, the combination of these two structures defines a non-trivial structure of Kähler fibration on $\curly{X}\to\curly{J}$, different from the one considered in Section 3.2.
In order to achieve our goal in this section, we will apply Theorem 3.2. As we have seen in Proposition 5.2, the integrability condition
@@
cuts a holomorphic submanifold $\curly{X}^{Higgs}$ of $(\curly{X},\mathbb{I})$. Given that $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$ is constant on $\curly{X}$, it provides a candidate for a coupling 2-form $\boldsymbol{\sigma}$ on $(\curly{X},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$, inducing the trivial Ehresmann connection on $\curly{X}\to\curly{J}$, regarded as the product $\curly{X}=\curly{J}\times{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$. However, one can readily check that the horizontal subspace of this connection is not tangent to the submanifold $\curly{X}^{Higgs}$, and hence is not appropriate for our goals. Motivated by this, we consider the Ehresmann connection $\Gamma^{\mathbb{I}}$ with horizontal subspace
$H^{\mathbb{I}}=\Big\{\left(\dot{J},0,-\tfrac{1}{2}\psi(J\dot{J})\right)\;\Big|\;\dot{J}\in T_{J}\curly{J}\Big\}.$
@@
where $\mathbf{I}$ denotes the complex structure along the fibres. Furthermore, its curvature $F_{\mathbb{I}}\in\Omega^{2}(\curly{X},V\curly{X})$ is of type $(1,1)$, and given explicitly by
@@
(5.6) $\begin{split}\widehat{\boldsymbol{\omega}}_{\mathbf{I}}(\Gamma^{\mathbb{I}}v_{1},\Gamma^{\mathbb{I}}v_{2})&=\int_{\Sigma}B(a_{1}\wedge a_{2})-\int_{\Sigma}B(\dot{\psi}_{1}\wedge\dot{\psi}_{2})\\ &-\frac{1}{2}\int_{\Sigma}B\left(\psi(J\dot{J}_{1})\wedge\dot{\psi}_{2}\right)-\frac{1}{2}\int_{\Sigma}B\left(\dot{\psi}_{1}\wedge\psi(J\dot{J}_{2})\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})\right).\end{split}$
Our next result proves that the Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler, by calculation of an explicit coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ via equation (3.3). Define a basic 2-form $\mu\in\Omega^{2}(\curly{X})$ by
@@
The Ehresmann connection $\Gamma^{\mathbb{I}}$ is Kähler. Furthermore, the following formula gives a coupling form for $\Gamma^{\mathbb{I}}$:
@@
$=-\widehat{\boldsymbol{\omega}}_{\mathbf{I}}((0,a,\dot{\psi}),F_{\mathbb{I}}(v_{1},v_{2})).\qed$
To finish this section, we provide a formula for the symmetric tensor on $\curly{X}$ associated to the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$, explicitly given by
@@
By construction, this coincides with the flat hyperkähler metric (2.8) along the fibres of $\curly{X}\to\curly{J}$. As in Corollary 3.7, $\mathbf{g}_{\mathbb{I}}$ is negative semi-definite along the horizontal directions of the connection $\Gamma^{\mathbb{I}}$.
@@
For any tangent vector $v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$ one has
(5.8) $\begin{split}\mathbf{g}_{\mathbb{I}}(v,v)&=\int_{\Sigma}B(a\wedge Ja)+\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{1}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}+\tfrac{1}{2}\psi(J\dot{J})\right)\right)\\ &-\frac{1}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right).\end{split}$
@@
Formula (5.8) follows from (5.7), using that $\Gamma^{\mathbb{I}}\circ\mathbb{I}=\mathbf{I}\circ\Gamma^{\mathbb{I}}$ (see Lemma 5.4). ∎
5.3. Hamiltonian action and the coupled Hitchin equations
We follow the notation of the previous section. Our aim now is to investigate a natural Hamiltonian action for the minimal couplings defined by the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$ associated to the connection $\Gamma^{\mathbb{I}}$, on the Kähler fibration $(\curly{X},\mathbb{I},\widehat{\boldsymbol{\omega}}_{\mathbf{I}})$ over the space of complex structures $\curly{J}$.
We fix a symplectic form $\omega$ on $\Sigma$, compatible with the orientation. Given $\alpha>0$ a real ‘coupling constant’ and $\varepsilon\in\{-1,1\}$, the family of minimal coupling symplectic structures of our interest is defined by
@@ 5.3. Hamiltonian action and the coupled Hitchin equations
The following result is the analogue of Lemma 3.8 in the present setup. Its proof is analogous, by application of Corollary 5.7 in this case, and is omitted.
@@
Let $v\in H^{\mathbb{I}}$ be a horizontal vector field at $(J,A,\psi)$, covering $\dot{J}\in T_{J}\curly{J}$. Then
@@
    If $\varepsilon=-1$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is negative definite along $H^{\mathbb{I}}$. In particular, $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,-1}$ is a non-degenerate symplectic structure.
@@
    If $\varepsilon=1$ and $\psi\neq 0$, $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ changes signature along the line $(J,A,\lambda\psi)\in\curly{X}$, for $\lambda\in\mathbb{R}$.
Our next goal is to prove that $(\curly{X},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ admits a Hamiltonian action by the extended gauge group $\widetilde{{\curly K}}$, determined by $\omega$ and the reduction $E_{K}\subset E_{G}$. For this, we will need some refined information about the coupling form $\boldsymbol{\sigma}^{\mathbb{I}}$. Consider the closed 2-form $\pi^{*}_{2}\omega_{\mathbf{I}}$ on $\curly{X}$, given by the pull-back of $\omega_{\mathbf{I}}=\operatorname{Re}\Omega_{\mathbf{J}}$ via the canonical projection $\pi_{2}\colon\curly{X}\to{\curly A}\times\Omega^{1}(\Sigma,E_{K}(\mathfrak{k}))$. Explicitly, for $v_{j}=(\dot{J}_{j},a_{j},\dot{\psi}_{j})\in T_{(J,A,\psi)}\curly{X}$, $j=1,2$, one has
@@
The 2-form $\pi^{*}_{2}\omega_{\mathbf{I}}$ restricts on the fibres to $\widehat{\boldsymbol{\omega}}_{\mathbf{I}}$, and hence, by the general theory of symplectic fibrations (see [GLS, Section 1.6]), one expects that there exists a 1-form $\lambda\in\Omega^{1}(\curly{X})$ such that
@@
Consider the 1-form $\lambda\in\Omega^{1}(\curly{X})$ defined by
@@
for $v=(\dot{J},a,\dot{\psi})\in T_{(J,A,\psi)}\curly{X}$. Then formula (5.13) holds. Furthermore, considering the $L^{2}$-norm of the ‘unitary Higgs field’,
@@
(5.14) $\boldsymbol{\sigma}^{\mathbb{I}}=\pi^{*}_{2}\omega_{{\curly A}}+\tfrac{i}{2}\partial_{\mathbb{I}}\bar{\partial}_{\mathbb{I}}\nu,$
@@
$=-4\pi^{*}_{2}\omega_{{\curly A}}+4\boldsymbol{\sigma}^{\mathbb{I}}.\qed$
Consider the extended gauge group $\widetilde{{\curly K}}$ of $h$ and $\omega$. As in [AGG1, Section 2.2], the group $\widetilde{{\curly K}}$ acts on $(\curly{X},\mathbb{I},\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon})$ preserving $\mathbb{I}$ and $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$, and covering the $\curly{H}$-action on $\curly{J}$ by push-forward. We are ready to prove the main result of this section. We use the same notation as in Proposition 3.9.
@@
$-\int_{\Sigma}f\left(\varepsilon S_{J}\omega+\alpha d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-\alpha d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right).$
@@
$=-\int_{\Sigma}fd^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right).\qed$
As in Proposition 3.9, the $\widetilde{{\curly K}}-$action produces a ‘coupling term’ on the base $\Sigma$ (cf. Remark 4.2), which interacts with the scalar curvature of the metric $g=\omega(,J)$. In particular, as in [AGG1], zeros of the moment map $\mu_{\widetilde{{\curly K}}}^{\mathbb{I}}$ are given by solutions to the coupled system of equations
(5.15) $\begin{split}F_{A}-\tfrac{1}{2}[\psi\wedge\psi]&=0,\\ \varepsilon S_{J}+\alpha\Lambda_{\omega}\left(d\left(B\left(\Lambda_{\omega}(d_{A}\psi)\right),\psi\right)-d^{c}\left(B\left(\psi,\Lambda_{\omega}(d_{A}(J\psi))\right)\right)\right)&=\varepsilon\frac{2\pi\chi(\Sigma)}{V}.\end{split}$
In the next section, we combine these equations with the integrability condition $\bar{\partial}_{A}\varphi=0$ for the Higgs field, in order to introduce the universal moduli space of Hitchin’s equations.